### v3

# Strictly convex norms and the local diameter two property

Abstract.
We introduce and study a strict monotonicity property of the norm in solid Banach lattices of real functions that prevents such spaces from having the local diameter two property. Then we show that any strictly convex 1-symmetric norm on $c_{0}(\Gamma)$ possesses this property. In the opposite direction, we show that any Banach space which is strictly convex renormable and contains a complemented copy of $c_{0}(\mathbb{N}),$ admits an equivalent strictly convex norm for which the space has the local diameter two property. In particular, this enables us to construct a strictly convex norm on $c_{0}(\Gamma),$ where $\Gamma$ is uncountable, for which the space has a 1-unconditional basis and the local diameter two property.
1. Introduction
A Banach space (or a norm $\|\cdot\|$ on a Banach space) $X$ is said to be strictly convex if for all $x,y\in X$ with $\frac{1}{2}\|x+y\|=\|x\|=\|y\|,$ we have $x=y.$ Another way of stating this is that the unit sphere $S_{X}$ of $X$ does not contain any non-trivial line segments.
A lot of effort has gone into understanding when a Banach space admits an equivalent strictly convex norm. It is known for example that every separable Banach space admits an equivalent strictly convex norm [Cla36, Theorem 9]. For non-separable spaces things get much more murky. Day showed that when $\Gamma$ is an uncountable set, $\ell_{\infty}(\Gamma)$ cannot be renormed to have a strictly convex norm, while $c_{0}(\Gamma)$ can [Day55]. Uncountability of $\Gamma$ is important here as $\ell_{\infty}(\mathbb{N})$ is strictly convex renormable (see [GMZ22, Remark 151 and Proposition 152]). Haydon [Hay90] showed that there exist uncountable scattered compacts $K$ for which $C(K)$ fails to be strictly convex renormable. At the present moment there does not exist an effective (checkable) criterion for when a (non-separable) Banach space admits an equivalent strictly convex norm or when it does not. For more on this topic the introduction in [OST12] is a good place to start.
Recall that a slice of the unit ball $B_{X}$ of a Banach space $X$ is a set of the form
$S(x^{*},\varepsilon):=\{x\in B_{X}:x^{*}(x)>1-\varepsilon\},$
where $x^{*}\in S_{X^{*}}$ and $\varepsilon>0.$ A Banach space $X$ is said to have the local diameter two property (LD2P) provided every slice of $B_{X}$ has diameter two. While the usual focus of renorming theory is to find an equivalent norm with the nicest properties possible our focus is different. We will study to what degree the LD2P and strict convexity can coexist. While at first sight the LD2P seems to be incompatible with strict convexity, there are examples showing that these properties indeed can coexist. In fact, even stronger properties are compatible with strict convexity. A Banach space is almost square if for every finite subset $x_{1},\ldots,x_{n}\in S_{X}$ and every $\varepsilon>0$ there exists $y\in S_{X}$ such that $\|x_{i}+y\|\leq 1+\varepsilon$ for all $i=1,\ldots,n$. Any almost square Banach space has the LD2P (see e.g [ALL16, Theorem 1.3 and Proposition 2.5]). Recall that a Banach space $X$ is M-embedded if $X^{***}=X^{*}\oplus_{1}X^{\perp}$, that is $X$ is an M-ideal in its bidual. Note that non-reflexive M-embedded spaces are almost square [ALL16, Corollary 4.3]. The quotient $C(\mathbb{T})/A$, where $C(\mathbb{T})$ is the space of continuous functions on the complex unit circle $\mathbb{T}$ and where $A$ is the disc algebra, is an example of a strictly convex Banach space with the LD2P. In fact, $C(\mathbb{T})/A$ is M-embedded and even has a smooth dual [HWW93, Remark IV.1.17]. Another example can be constructed as follows: Let $\varphi$ be the function on $c_{0}(\mathbb{N})$ defined by $\varphi(x)=\sum_{n=1}^{\infty}x^{2n}_{n}$ where $x=(x_{n}).$ The Nakano norm on $c_{0}(\mathbb{N})$ is defined by
$\|x\|=\inf\{\lambda>0:\varphi(x/\lambda)\leq 1\}$
for every $x\in c_{0}(\mathbb{N})$. The space $c_{0}(\mathbb{N})$ with the Nakano norm is strictly convex and is almost square [AHN+16]. A third and recent example is the strictly convex renorming of $L_{1}[0,1]$ constructed in [NPoTV24]. This example is not almost square, but has the LD2P.
While the LD2P and strict convexity can coexist in a Banach space $X$, strict convexity in $X^{**}$ is asking too much, at least if we ask $X$ to have the property that all finite convex combinations of slices of its unit ball have diameter two. (Whether the same conclusion holds only assuming the LD2P, is an open question). Indeed, in this case not only does $X^{**}$ fail to be strictly convex, it also fails to be smooth [ALNT16, Corollary 2.6]. Another result in this direction is that if $X$ has a bimonotone basis and the property that all non-empty relatively weakly open subsets of its unit ball have diameter two, then $X^{**}$ is not strictly convex [ALNT16, Proposition 2.10]. These results point to the fact that constructing strictly convex Banach spaces with the LD2P is not completely straightforward.
Note that all the examples above with the LD2P and strictly convex norms, are separable. If $\Gamma$ is an uncountable set we know that Day’s norm on $c_{0}(\Gamma)$ is strictly convex (and even locally uniformly rotund), but it is not immediately obvious how to give $c_{0}(\Gamma)$ a strictly convex norm with the LD2P. If the norm of a Banach space $X$ is locally uniformly rotund then every $x\in S_{X}$ is strongly exposed, so the LD2P and local uniformly rotundity cannot coexist. In the present paper, we construct the first, to the best of our knowledge, example of a strictly convex renorming of $c_{0}(\Gamma)$ with the LD2P.
The rest of the paper is organized as follows: After some preliminary results in Section 2, we introduce in Section 3 a strict monotonicity property of the norm in solid Banach lattices of real functions and show that this property cannot coexist with the LD2P.
In Section 4 we show that under some assumptions on the norm, solid Banach lattices of real functions possess the strict monotonicity property just mentioned. In particular, we show that for any set $\Gamma,$ any strictly convex norm on $c_{0}(\Gamma)$ with a 1-symmetric basis, fails the LD2P.
Finally, in Section 5 we prove that every strictly convex renormable Banach space which contains a complemented copy of $c_{0}(\mathbb{N})$ has an equivalent norm which is both strictly convex and almost square. In particular, this enables us to prove that for any infinite set $\Gamma$, $c_{0}(\Gamma)$ has a strictly convex renorming with a 1-unconditional basis that is almost square.
2. Preliminaries
Let us start by recalling some definitions and results that we will need later in the paper.
A solid Banach lattice of real functions on some set $\Gamma$ is a Banach space $(X,\|\cdot\|)$ consisting of real-valued bounded functions $x:\Gamma\to\mathbb{R}$ with the following properties
  • if $x\in X$ and $y:\Gamma\to\mathbb{R}$ with $|y(\gamma)|\leq|x(\gamma)|$ for all $\gamma\in\Gamma$, then $y\in X$ and $\|y\|\leq\|x\|$;
  • for every $\alpha\in\Gamma,$ the function $e_{\alpha}:=\delta_{\alpha}$, where $\delta_{\alpha}(\gamma)=1$ if $\gamma=\alpha,$ and $\delta_{\alpha}(\gamma)=0$ if $\gamma\neq\alpha$, belongs to $X$.
We will assume that $\|e_{\gamma}\|=1$ for every $\gamma\in\Gamma.$ Hence $\ell_{1}(\Gamma)\subseteq X\subseteq\ell_{\infty}(\Gamma)$ and $\|\cdot\|_{\infty}\leq\|\cdot\|\leq\|\cdot\|_{1}$. When $\Gamma$ is the set $\mathbb{N}$ of natural numbers, we obtain the well known class of Köthe sequence spaces.
For $\gamma\in\Gamma,$ we let $e_{\gamma}^{*}\in X^{*}$ denote the biorthogonal functional to $e_{\gamma}.$ Note that for $x\in X$ we have $e_{\gamma}^{*}(x)=x(\gamma).$
Clearly $(e_{\gamma})_{\gamma\in\Gamma}$ is a 1-unconditional basic set in $X,$ i.e. a 1-unconditional basis for the norm closure of the linear span of $\{e_{\gamma}:\gamma\in\Gamma\}$ in $X.$ Hence if $c_{00}(\Gamma)$, the space of finitely supported real-valued functions on $\Gamma,$ is dense in $X$, the unit vectors $(e_{\gamma})_{\gamma\in\Gamma}$ form a 1-unconditional basis for $X$. Recall that if $\Gamma$ is a non-empty set, then $(e_{\gamma})_{\gamma\in\Gamma}$ is called an unconditional basis for a Banach space $X$ if for every $x\in X$ there is a unique family of real numbers $(a_{\gamma})_{\gamma\in\Gamma}$ such that $x=\sum_{\gamma\in\Gamma}a_{\gamma}e_{\gamma}$ in the sense that for every $\varepsilon>0$ there is a finite set $F\subset\Gamma$ such that $\|x-\sum_{\gamma\in G}a_{\gamma}e_{\gamma}\|<\varepsilon$ for every $G\supset F.$ If moreover, for any finite set $F$ in $\Gamma$ and any set of real numbers $(a_{\gamma})_{\gamma\in F}$ and $(b_{\gamma})_{\gamma\in F}$ with $|b_{\gamma}|\leq|\alpha_{\gamma}|$ for $\gamma\in F,$ we have
$\left\|\sum_{\gamma\in F}b_{\gamma}e_{\gamma}\right\|\leq\left\|\sum_{\gamma\in F}a_{\gamma}e_{\gamma}\right\|,$
then we say that $(e_{\gamma})_{\gamma\in\Gamma}$ is 1-unconditional.
If $X$ is a solid Banach lattice of real functions on $\Gamma,$ we say that $X$ is 1-symmetric if for every $x=x(\gamma)\in X$ and every permutation $\pi:\Gamma\to\Gamma$ we have $x_{\pi}=x(\pi(\gamma))\in X$ and $\|x_{\pi}\|=\|x\|$. We say that the norm of the lattice $X$ is strictly monotone if $\|y\|<\|x\|$ whenever $|y|<|x|$, that is if $|y(\gamma)|\leq|x(\gamma)|$ for all $\gamma\in\Gamma$ and $|y(\beta)|<|x(\beta)|$ for some $\beta\in\Gamma$.
Let $(e_{\gamma})_{\gamma\in\Gamma}$ be a 1-unconditional basis for a Banach space $X.$ For each $\gamma\in\Gamma$ let $e_{\gamma}^{*}\in X^{*}$ denote the biorthogonal functional of $e_{\gamma}.$ We say that $(e_{\gamma},e_{\gamma}^{*})_{\gamma\in\Gamma}$ is shrinking if $\operatorname{\overline{span}}\{e_{\gamma}^{*}:\gamma\in\Gamma\}=X^{*}.$
The following results will be used several times in the paper. The first one is James’ classical representation of the bidual of a Banach space with a shrinking basis in case the shrinking basis is uncountable and unconditional. We include a proof for easy reference.
Proposition 2.1.
Let $X$ be a Banach space with a shrinking 1-unconditional basis $(e_{\gamma},e_{\gamma}^{*})_{\gamma\in\Gamma}$, and $\mathcal{F}$ the set $\{F\subset\Gamma:|F|<\infty\}$ ordered by inclusion. Then for every $x^{**}\in X^{**}$ we have
$\|x^{**}\|=\sup_{F\in\mathcal{F}}\left\|\sum_{\gamma\in F}x^{**}(e_{\gamma}^{*})e_{\gamma}\right\|=\lim_{\mathcal{F}}\left\|\sum_{F\in\mathcal{F}}x^{**}(e_{\gamma}^{*})e_{\gamma}\right\|.$
Proof.
Let $F\subset\Gamma$ be finite and $P_{F}:X\to X$ the projection onto $\operatorname{span}\{e_{\gamma}:\gamma\in F\}$ defined by
$P_{F}(x)=\sum_{\gamma\in F}e_{\gamma}^{*}(x)e_{\gamma}.$
Since the basis is shrinking, we have
(1) $P_{F}^{**}(x^{**})=\sum_{\gamma\in F}x^{**}(e_{\gamma}^{*})e_{\gamma},$
for every $x^{**}\in X^{**}.$ Indeed if $\gamma_{0}\in\Gamma$ and $x^{*}=e_{\gamma_{0}}^{*},$ then
(2) $\langle P_{F}^{**}(x^{**}),x^{*}\rangle=\langle x^{**},P_{F}^{*}(e_{\gamma_{0}}^{*})\rangle=\begin{cases}x^{**}(e_{\gamma_{0}}^{*}),&\gamma_{0}\in F\\ 0,&\gamma_{0}\notin F,\end{cases}$
so
$\langle P_{F}^{**}(x^{**}),x^{*}\rangle=\left\langle\sum_{\gamma\in F}x^{**}(e_{\gamma}^{*})e_{\gamma},x^{*}\right\rangle.$
From this in tandem with linearity and continuity of the functionals involved and the fact that $\operatorname{\overline{span}}\{e_{\gamma}^{*}:\gamma\in\Gamma\}=X^{*},$ the equality (1) follows. Moreover, since the basis is 1-unconditional,
$\|x^{**}\|=\sup_{F\in\mathcal{F}}\|P_{F}^{**}x^{**}\|.$
Indeed, because of (2) and since $\operatorname{\overline{span}}\{e_{\gamma}^{*}:\gamma\in\Gamma\}=X^{*},$ the net $(P_{F}^{**}x^{**})_{F\in\mathcal{F}},$ converges weak to $x^{**}.$ Thus by weak lower semi-continuity of the norm we have $\|x^{**}\|\leq\liminf_{\mathcal{F}}\|P_{F}^{**}x^{**}\|\leq\sup_{F\in\mathcal{F}}\|P_{F}^{**}x^{**}\|.$ Also, since $(e_{\gamma},e_{\gamma}^{*})_{\gamma\in\Gamma}$ is 1-unconditional (in particular monotone) we have $\|P_{F}\|=1,$ so we can conclude
$\|x^{**}\|=\sup_{F\in\mathcal{F}}\|P_{F}^{**}x^{**}\|=\lim_{\mathcal{F}}\|P_{F}^{**}x^{**}\|.\qed$
The next result is well known. For a proof see that of Theorem 2.6 (ii) implies (iii) in [MOTZ07].
Proposition 2.2.
Let $X$ be a solid Banach lattice of real functions. If the norm of $X$ is strictly convex, then $X$ is strictly monotone.
The converse does not hold in general, just consider $X=\ell_{1}$. But if $(X,\|\cdot\|)$ is a solid Banach lattice of real functions on some set $\Gamma$ such that $\|\cdot\|_{\infty}\leq\|\cdot\|,$ we have from [MOTZ07, Theorem 2.6], that the following statements are equivalent:
  1. (i)
    $X$ admits a pointwise lower semi-continuous strictly convex norm;
  2. (ii)
    $X$ admits a lattice pointwise lower semi-continuous strictly convex norm;
  3. (iii)
    $X$ admits a pointwise lower semi-continuous strictly monotone norm (in [MOTZ07] the term strictly lattice norm is used).
3. A uniform strict monotonicity property of the norm
Throughout this section $X$ will be a solid Banach lattice of real functions on a set $\Gamma$ with a (normalized) 1-unconditional basic set $(e_{\gamma})_{\gamma\in\Gamma}.$ In this section we will introduce a property that will prevent $X$ from having the LD2P. The Nakano norm on $c_{0}$ shows that a strictly monotone norm is not enough to prevent the LD2P. On the other hand, the space $c_{0}\oplus_{1}\mathbb{R}$ has a 1-unconditional basis, fails the LD2P, and the norm is not strictly monotone. The last example can serve as a motivation for our property.
Definition 3.1.
For each $\alpha\in\Gamma$ and $\varepsilon>0$ define
$E_{\alpha}(\varepsilon):=\sup\left\{\|x-e_{\alpha}^{*}(x)e_{\alpha}\|:x\in S_{X},e_{\alpha}^{*}(x)>1-\varepsilon\right\}.$
If there exist $\alpha\in\Gamma$ and $\varepsilon>0$ such that $E_{\alpha}(\varepsilon)<1$, then we say that $X$ is uniformly strictly monotone at $\alpha\in\Gamma$.
We note that $E_{\alpha}$ is a non-decreasing function of $\varepsilon$ since we are taking supremum over a smaller set.
Proposition 3.2.
If there exists $\alpha\in\Gamma$ such that $X$ is uniformly strictly monotone at $\alpha$, then $X$ fails to have the LD2P.
Proof.
Assume $X$ is uniformly strictly monotone at $\alpha\in\Gamma.$ Let $\varepsilon>0$ such that $E_{\alpha}(\varepsilon)<1$. We may and will assume that $\varepsilon<2/3$. We will show that the diameter of the slice $S(e_{\alpha}^{*},\varepsilon)$ is bounded away from $2$.
Claim 3.3.
There exists a $C>0$ such that
$\left\|x-\left(e_{\alpha}^{*}(x)-\frac{\varepsilon}{2}\right)e_{\alpha}\right\|\leq 1-C$
for all $x\in S(e_{\alpha}^{*},\varepsilon)$.
Proof of claim.
Let $C:=(1-3\varepsilon/2)(1-E_{\alpha}(\varepsilon))>0$.
If $x\in S(e_{\alpha}^{*},\varepsilon)$ we set
$\lambda=\frac{e_{\alpha}^{*}(x)-\varepsilon/2}{e_{\alpha}^{*}(x)}.$
Then, since $e_{\alpha}^{*}(x)\leq 1$ and $\varepsilon<2/3$,
$1>\lambda>\frac{1-\varepsilon-\varepsilon/2}{e_{\alpha}^{*}(x)}>1-\frac{3}{2}\varepsilon>0$
and
$x-\left(e_{\alpha}^{*}(x)-\frac{\varepsilon}{2}\right)e_{\alpha}=(1-\lambda)x+\lambda(x-e_{\alpha}^{*}(x)e_{\alpha}),$
so by convexity of the norm
$\left\|x-\left(e_{\alpha}^{*}(x)-\frac{\varepsilon}{2}\right)e_{\alpha}\right\|$ $\leq(1-\lambda)\|x\|+\lambda\|x-e_{\alpha}^{*}(x)e_{\alpha}\|$
$\leq 1-\lambda+\lambda E_{\alpha}(\varepsilon)\leq 1-C$
as claimed. ∎
Let $y,z\in S(e_{\alpha}^{*},\varepsilon)$. We use unconditionality, i.e. $\|u\|\leq\|v\|$ whenever $|u|\leq|v|$, and the above claim and get
$\|y-z\|$ $=\left\|e_{\alpha}^{*}(y-z)e_{\alpha}+\bigl{(}y-e_{\alpha}^{*}(y)e_{\alpha}\bigr{)}-\bigl{(}z-e_{\alpha}^{*}(z)e_{\alpha}\bigr{)}\right\|$
$\leq\left\|\varepsilon e_{\alpha}+\bigl{(}y-e_{\alpha}^{*}(y)e_{\alpha}\bigr{)}-\bigl{(}z-e_{\alpha}^{*}(z)e_{\alpha}\bigr{)}\right\|$
$\leq\left\|\bigl{(}y-e_{\alpha}^{*}(y)e_{\alpha}\bigr{)}+\frac{\varepsilon}{2}e_{\alpha}\right\|+\left\|\bigl{(}z-e_{\alpha}^{*}(z)e_{\alpha}\bigr{)}-\frac{\varepsilon}{2}e_{\alpha}\right\|$
$=\left\|y-\left(e_{\alpha}^{*}(y)-\frac{\varepsilon}{2}\right)e_{\alpha}\right\|+\left\|z-\left(e_{\alpha}^{*}(z)-\frac{\varepsilon}{2}\right)e_{\alpha}\right\|$
$\leq 2(1-C).$
Hence the diameter of $S(e_{\alpha}^{*},\varepsilon)$ is strictly less than $2$. ∎
4. Spaces with 1-unconditional bases that have uniformly strictly monotone norms
Let $X$ be a solid Banach lattice of real functions on a set $\Gamma$ with a (normalized) 1-unconditional basic set $(e_{\gamma})_{\gamma\in\Gamma}.$ For $A\subset\Gamma,$ let $\mathbbm{1}_{A}\in\ell_{\infty}(\Gamma)$ denote the characteristic function on $A.$ Note that $\mathbbm{1}_{A}\in X$ if $A$ is finite. We define
$\mathfrak{L_{F}}=\sup\left\{\left\|\mathbbm{1}_{A}\right\|:A\subset\Gamma,|A|<\infty\right\}\text{ and }\mathfrak{L}=\left\|\mathbbm{1}_{\Gamma}\right\|.$
Clearly $\mathfrak{L}\geq\mathfrak{L_{F}}.$ Also, since $(e_{\gamma})_{\gamma\in\Gamma}$ is normalized we have $\mathfrak{L_{F}}\geq 1$, and if $X$ is strictly monotone we have $\mathfrak{L_{F}}>1$. When we write $\mathfrak{L}<\infty,$ we implicitly assume that $\mathbbm{1}_{\Gamma}$ is an element in $X.$ Clearly $\|\mathbbm{1}_{\Gamma}\|<\infty$ is equivalent to $X$ being lattice isomorphic to $\ell_{\infty}(\Gamma).$
For a finite subset $F$ of $\Gamma$ let $P_{F}:X\to X$ be the natural projection onto the set $\operatorname{span}\{e_{\gamma}:\gamma\in F\}.$
If $F\subset\Gamma$ is finite and $(a_{\gamma})_{\gamma\in F}$ is a sequence of scalars, then
$\max_{\gamma\in F}|a_{\gamma}|\leq\left\|\sum_{\gamma\in F}a_{\gamma}e_{\gamma}\right\|\leq\max_{\gamma\in F}|a_{\gamma}|\cdot\left\|\sum_{\gamma\in F}e_{\gamma}\right\|\leq\mathfrak{L_{F}}\cdot\max_{\gamma\in F}|a_{\gamma}|.$
Thus $\operatorname{\overline{span}}\{e_{\gamma}:\gamma\in\Gamma\}$ is isomorphic to $c_{0}(\Gamma)$ if and only if $\mathfrak{L_{F}}<\infty$.
Let $X$ be $c_{0}(\mathbb{N})$ with the Nakano norm. The canonical basis for $c_{0}(\mathbb{N})$, $(e_{n})_{n=1}^{\infty}$, is then a normalized strictly monotone 1-unconditional basis for $X$. If we want to show that $X$ has the LD2P, then we can make use of the following fact: For every $j\in\mathbb{N}$ we have that for any fixed $r\in(0,1)$
$\inf_{n\neq j}\left\{\|e_{j}+re_{n}\|\right\}=1.$
In particular, this holds for $r=1/\mathfrak{L_{F}}$. Actually, we can replace $e_{j}$ with any $x\in S_{X}$ with finite support in the infimum above. Hence inside every slice $S(x^{*},\varepsilon)$ of $B_{X}$ we can for every $r\in(0,1)$ find elements of the form $x\pm re_{n}.$ This tells us that every such slice has diameter two. We will return to this example in Section 5.
Theorem 4.1.
Let $X$ be a strictly monotone solid Banach lattice of real functions on the set $\Gamma$ with a basic set $(e_{\gamma})_{\gamma\in\Gamma}.$
  1. (i)
    Assume $\mathfrak{L_{F}}<\infty$ and for every $x\in X$ we have $\|x\|=\sup_{F\subset\Gamma,|F|<\infty}\|P_{F}x\|.$ If there exists $\alpha\in\Gamma$ such that
    $\inf_{\beta\neq\alpha}\left\{\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\beta}\right\|\right\}>1,$
    then $X$ fails the LD2P. In particular, any equivalent renorming of $c_{0}(\Gamma)$ such that the canonical basis $(e_{\gamma})_{\gamma\in\Gamma}$ is strictly monotone and normalized, and there exists $\alpha\in\Gamma$ with
    $\inf_{\beta\neq\alpha}\left\{\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\beta}\right\|\right\}>1,$
    fails the LD2P.
  2. (ii)
    Assume that $X$ is lattice isomorphic to $\ell_{\infty}(\mathbb{N}).$ If there exists $k\in\mathbb{N}$ such that
    $\inf_{n\neq k}\left\{\left\|e_{k}+\frac{1}{\mathfrak{L}}e_{n}\right\|\right\}>1,$
    then $X$ fails the LD2P.
Remark 4.2.
Note that it follows from Day’s proof in [Day55] that when $\Gamma$ is uncountable, $\ell_{\infty}(\Gamma)$ does not have an equivalent strictly monotone norm.
Proof.
Let $\alpha\in\Gamma$. We aim to show that $X$ is uniformly strictly monotone at $\alpha$ and the proof of (i) and (ii) are the same up to a point.
Assume $K>1$ is a real number such that
$l:=\inf_{\beta\neq\alpha}\left\{\left\|e_{\alpha}+\frac{1}{K}e_{\beta}\right\|\right\}>1.$
Fix $\eta<1$ such that
$\eta>\frac{K+1}{lK+1}.$
Note that since $l\leq 2$
$\eta>\frac{K+1}{lK+1}>\frac{K+1}{l(K+1)}=\frac{1}{l}\geq\frac{1}{2}.$
Claim 4.3.
There exists $\mathfrak{a}<1/K$ such that for all $\eta e_{\alpha}+ae_{\beta}\in B_{X}$ with $a\in\mathbb{R}$ and $\beta\neq\alpha,$ we have $|a|\leq\mathfrak{a}$.
Proof of claim.
For notational reasons, let us introduce the function
$f(\tau):=\tau\eta l-(1-\tau)\eta-(1-\eta)\tau/K.$
Rewriting the assumption
$\eta>\frac{K+1}{lK+1}$
gives $f(1)=\eta l-(1-\eta)/K>1$. Since $f$ is continuous we can find $\theta\in(0,1)$ such that
$f(\theta)=\theta\eta l-(1-\theta)\eta-(1-\eta)\theta/K>1.$
Set $\mathfrak{a}:=\theta/K.$ Let $\|\eta e_{\alpha}+ae_{\beta}\|\leq 1$ for some $\beta\in\Gamma\setminus\{\alpha\}$ and $a\in\mathbb{R}.$ Define
$g_{\beta}(t):=\left\|\eta e_{\alpha}+te_{\beta}\right\|.$
Since
$\eta e_{\alpha}+\theta e_{\beta}/K=\theta\eta(e_{\alpha}+e_{\beta}/K)+(1-\theta)\eta e_{\alpha}+(1-\eta)\theta e_{\beta}/K,$
we have
$g_{\beta}(\theta/K)$ $=\|\eta e_{\alpha}+\theta e_{\beta}/K\|$
$\geq\theta\eta\|e_{\alpha}+e_{\beta}/K\|-(1-\theta)\eta\|e_{\alpha}\|-(1-\eta)\theta/K\|e_{\beta}\|$
$\geq\theta\eta l-(1-\theta)\eta-(1-\eta)\theta/K$
$=f(\theta)$
$>1\geq g_{\beta}(a).$
By 1-unconditionality of the basic set $(e_{\gamma})_{\gamma\in\Gamma},$ we have that $g_{\beta}$ is even and is non-decreasing for $t\geq 0.$ Hence $|a|\leq\theta/K=\mathfrak{a}.$
We are now ready to prove that $X$ is uniformly strictly monotone at $\alpha$. Let $0<\varepsilon<1-\eta$ be such that $\mathfrak{a}K<1-2\varepsilon$. We will show that
$E_{\alpha}(\varepsilon)=\sup\{\|x-e_{\alpha}^{*}(x)e_{\alpha}\|:x\in S_{X},e_{\alpha}^{*}(x)>1-\varepsilon\}\leq 1-\varepsilon.$
Let $x\in S_{X}$ and assume $e_{\alpha}^{*}(x)>1-\varepsilon$. Let $\beta\neq\alpha$. Now for some $z_{\beta}\in B_{X}$ we have
$x=e_{\alpha}^{*}(x)e_{\alpha}+e_{\beta}^{*}(x)e_{\beta}+z_{\beta}$
and by strict monotonicity and $1$-unconditionality
$1=\|x\|\geq\|e_{\alpha}^{*}(x)e_{\alpha}+e_{\beta}^{*}(x)e_{\beta}+z_{\beta}\|>\|\eta e_{\alpha}+e_{\beta}^{*}(x)e_{\beta}\|.$
From Claim 4.3 we get that $|e_{\beta}^{*}(x)|\leq\mathfrak{a}<1/K$ for all $\beta\neq\alpha$.
(i). We assume $\mathfrak{L_{F}}<\infty.$ Put $K:=\mathfrak{L_{F}},$ and note that $\mathfrak{L_{F}}>1$ since $X$ is strictly monotone. Therefore $K>1,$ and we get by the assumption that there exists a finite subset $F\subset\Gamma$ such that
$\|x-e_{\alpha}^{*}(x)e_{\alpha}\|$ $\leq\left\|\sum_{\beta\in F\setminus\{\alpha\}}e_{\beta}^{*}(x)e_{\beta}\right\|+\varepsilon\leq\mathfrak{a}\left\|\sum_{\beta\in F\setminus\{\alpha\}}e_{\beta}\right\|+\varepsilon$
$\leq\mathfrak{a}\left\|\sum_{\beta\in F\cup\{\alpha\}}e_{\beta}\right\|+\varepsilon<\mathfrak{a}\mathfrak{L_{F}}+\varepsilon=1-\varepsilon.$
This proves that $X$ is uniformly strictly monotone at $\alpha,$ and we have from Proposition 3.2 that $X$ fails the LD2P.
For the particular case, let $X$ be an equivalent renorming of $c_{0}(\Gamma)$ such that the canonical basis $(e_{\gamma})_{\gamma\in\Gamma}$ is strictly monotone and normalized. We then have that $X$ is a strictly monotone solid Banach lattice consisting of functions $x:\Gamma\to\mathbb{R}.$ Now we only need to note that under the assumptions on $X$ we have $\mathfrak{L_{F}}<\infty$ and from Proposition 2.1 that $\|x\|=\sup_{F\subset\Gamma,|F|<\infty}\|P_{F}x\|$ for every $x\in X.$
(ii). We assume $\Gamma=\mathbb{N}$ and $\mathfrak{L}<\infty.$ Put $K:=\mathfrak{L}$ and note that $\mathfrak{L}>1$ since $X$ is strictly monotone. Therefore $K>1,$ and with $k=\alpha$ we get
$\|x-e_{k}^{*}(x)e_{k}\|\leq\left\|\mathfrak{a}\mathbbm{1}_{\operatorname{supp}(x)\setminus\{k\}}\right\|<\mathfrak{a}\left\|\mathbbm{1}_{\Gamma}\right\|=\mathfrak{a}\mathfrak{L}<1-\varepsilon,$
so $E_{k}(\varepsilon)<1-\varepsilon$ and $X$ is therefore uniformly strictly monotone at $k.$ As in (i) we get that $X$ fails the LD2P by Proposition 3.2. ∎
Theorem 4.4.
Let $\Gamma$ be a non-empty set and let $\|\cdot\|$ be an equivalent norm on $c_{0}(\Gamma)$ such that the canonical basis $(e_{\gamma})_{\gamma\in\Gamma}$ is normalized and 1-symmetric. If $\|\cdot\|$ is strictly monotone, in particular if $\|\cdot\|$ strictly convex, then it fails the LD2P.
Proof.
Let $X$ be a renorming of $c_{0}(\Gamma)$ such that the canonical basis $(e_{\gamma})_{\gamma\in\Gamma}$ is strictly monotone, normalized, and 1-symmetric. Choose any $\alpha\in\Gamma$. By Theorem 4.1 (i), it is enough to show that
$\inf_{\beta\neq\alpha}\left\{\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\beta}\right\|\right\}>1.$
But by symmetry we have
$\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\beta}\right\|=\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\gamma}\right\|$
for all $\beta,\gamma\in\Gamma\setminus\{\alpha\}$, so it is enough to observe that for $\beta\neq\alpha$
$\left\|e_{\alpha}+\frac{1}{\mathfrak{L_{F}}}e_{\beta}\right\|>\|e_{\alpha}\|=1,$
by strict monotonicity.
The strictly convex case now follows from Proposition 2.2. ∎
Remark 4.5.
It was shown in [HZ95, Corollary 6] that $\ell_{\infty}(\mathbb{N})$ admits no equivalent strictly convex 1-symmetric norm. It can also be shown that $\ell_{\infty}(\mathbb{N})$ does not admit a strictly monotone 1-symmetric norm.
Question 4.6.
Let $X$ be a strictly monotone renorming of $c_{0}(\mathbb{N})$ such that the canonical basis $(e_{n})_{n=1}^{\infty}$ is normalized. Are then the following two conditions equivalent?
  1. (i)
    $X$ fails the LD2P;
  2. (ii)
    There exists $j\in\mathbb{N}$ such that $\inf_{n\neq j}\left\{\left\|e_{j}+\frac{1}{\mathfrak{L_{F}}}e_{n}\right\|\right\}>1.$
Let us end this section with some remarks about the assumption $\mathfrak{L}_{\mathfrak{F}}<\infty$ in Theorem 4.1.
First let us give an example of a Banach space $X$ with a 1-symmetric basis and a strictly convex norm, which has the LD2P, but which is not uniformly strictly monotone at any coordinate and with $\mathfrak{L}_{\mathfrak{F}}=\infty$.
Example 4.7.
Define a function $M:[0,\infty)\to[0,\infty)$ by
$M(t)=\begin{cases}0,&t=0;\\ \frac{1}{4}e^{2}\cdot e^{-1/t},&t\in(0,\frac{1}{2});\\ t^{2},&t\geq\frac{1}{2}.\end{cases}$
Then $M$ is a convex, non-decreasing continuous function satisfying $M(0)=0$ and $\lim_{t\to\infty}M(t)=\infty$. If we consider the space $\ell_{M}$ of all sequences $x=(a_{1},a_{2},\ldots)$ of scalars such that $\Phi(x/\lambda):=\sum_{n=1}^{\infty}M(|a_{n}|/\lambda)<\infty$ for some $\lambda>0$, then $\ell_{M}$ is a Banach space with norm
$\|x\|=\inf\{\lambda>0:\sum_{n=1}^{\infty}M(|a_{n}|/\lambda)\leq 1\}.$
The unit vectors $\{e_{n}\}_{n=1}^{\infty}$ form a 1-symmetric basis of $h_{M}=\operatorname{\overline{span}}\{e_{n}\}_{n=1}^{\infty}$ (see e.g. [LT77, p. 115]).
Let us show that $X=h_{M}$ is also strictly convex, has the LD2P (it is actually M-embedded), is not uniformly strictly monotone at any coordinate, but $\mathfrak{L}_{\mathfrak{F}}=\infty$.
A straightforward computation shows that $M$ is strictly convex, so $h_{M}$ is strictly convex (see e.g. [Che96, p. 56]).
From Example III.1.4 in [HWW93] we have that $h_{M}$ is M-embedded since the complementary Orlicz function $M^{*}$ satisfies the $\Delta_{2}$ condition at zero while $M$ fails it. In particular, $h_{M}$ has the LD2P by [BGLPPRP04, Proposition 2.1]. Let us prove the claimed properties of $M$ and $M^{*}$. We have that $M^{\prime}(t)=M(t)\frac{1}{t^{2}}$ on $(0,\frac{1}{2})$ and hence
$\lim_{t\to 0^{+}}\frac{tM^{\prime}(t)}{M(t)}=\lim_{t\to 0^{+}}\frac{tM(t)/t^{2}}{M(t)}=\lim_{t\to 0^{+}}\frac{1}{t}=+\infty.$
Since the above limit is infinite $M$ fails the $\Delta_{2}$-condition at zero (see e.g. [LT77, p. 140]). We have that $M^{\prime}(t)$ is continuous and strictly increasing on $0,\infty)$ and has a continuous and strictly increasing inverse $q(t)$. The complementary function to $M$ is given by $M^{*}(u)=\int_{0}^{u}q(t)\,dt$ and $M^{*}$ satisfies the $\Delta_{2}$-condition at zero (see the remarks following Proposition 4.b.2 in [LT77, p. 148]) since the above limit is strictly greater than $1$ (here we use that $q$ is continuous).
Next let us show that $\mathfrak{L}_{\mathfrak{F}}=\infty$. For $s_{n}=\sum_{i=1}^{n}e_{i}$ in $h_{M}$ we have
$\Phi(s_{n}/\lambda)=\sum_{i=1}^{n}M(1/\lambda)=nM(1/\lambda).$
It is well known that the infimum in the norm is attained so we need only solve $\Phi(s_{n}/\lambda)=1$ to get $M(1/\lambda)=1/n$. For $n\geq 4$ we use $M(1/\lambda)=\frac{1}{4}e^{2}e^{-\lambda}$ and get
(3) $\|s_{n}\|=\left\|\sum_{i=1}^{n}e_{i}\right\|=\lambda=2-2\ln(2)+\ln(n)$
From this we see that $\mathfrak{L}_{\mathfrak{F}}=\infty$.
Finally, we show that $h_{M}$ is not uniformly strictly monotone at any coordinate. By 1-symmetry of the basis we need only consider the first coordinate.
Let $0<d<1/2$. For $n\in\mathbb{N}$ define
$x_{n}=(1-d)e_{1}+\sum_{i=1}^{n}\frac{1}{k}e_{i+1}$
where $k$ is chosen (depending on $n$) such that $\|x_{n}\|=1$. Using (3) we have
$\|x_{n}-e_{1}^{*}(x_{n})e_{1}\|=\frac{1}{k}\left\|\sum_{i=1}^{n}e_{i+1}\right\|=\frac{\|s_{n}\|}{k}=\frac{2-2\ln(2)+\ln(n)}{k}.$
Let us find $k$. Solve
$\Phi(x_{n})=(1-d)^{2}+\frac{n}{4}e^{2}e^{-k}=1$
for $k$ to get
$e^{-k}=e^{-2}\frac{4}{n}(2d-d^{2})$
hence
$k=2-2\ln(2)+\ln(n)-\ln(2d-d^{2}).$
We therefore have
$\|x_{n}-e_{1}^{*}(x_{n})e_{1}\|$ $=\frac{2-2\ln(2)+\ln(n)}{2-2\ln(2)+\ln(n)-\ln(2d-d^{2})}$
$=1+\frac{\ln(2d-d^{2})}{2-2\ln(2)+\ln(n)-\ln(2d-d^{2})}.$
Since $d$ is fixed and $0<d<1/2$ we have that $\ln(2d-d^{2})$ is a fixed negative number and $\ln(n)$ can be made arbitrarily large, hence $\lim_{n\to\infty}\|x_{n}-e_{1}^{*}(x_{n})e_{1}\|=1$. It follows that $h_{M}$ is not uniformly strictly monotone at any coordinate.
As proved in [HWW93, III.Examples 1.4], if $M$ is an Orlicz function which fails the $\Delta_{2}$-condition at zero, while its complementary function $M^{*}$ satisfies it, then the Orlicz sequence space $h_{M}$ is M-embedded and thus has the LD2P. The same conclusion holds for the preduals $d(w,1)_{*}$ of the Lorentz sequence space $d(w,1)$. Example 4.7 shows that there exist strictly convex Orlicz sequence spaces $h_{M}$ with a 1-symmetric basis and the LD2P. As for Lorentz sequences spaces $d(w,1),$ however, their preduals $d(w,1)_{*}$ are never strictly convex since their canonical basis is monotone, but not strictly monotone. We will come back to these preduals once more in the next section.
Remark 4.8.
In general, $\mathfrak{L_{F}}=\infty$ does not tell us anything about uniform strict monotonicity at a coordinate or the LD2P.
Any Orlicz space with $M$ non-degenerate (hence $M$ strictly increasing) satisfies $\mathfrak{L_{F}}=\infty$. The Orlicz space $h_{M}$ in Example 4.7 has the LD2P and is not uniform strict monotonicity at any coordinate. On the other hand, $\ell_{p}$, $1\leq p<\infty$, clearly does not have the LD2P. If $x\in B_{\ell_{p}}$ with $e_{1}^{*}(x)=(1-d)$ we have $\|x-e_{1}^{*}(x)e_{1}\|=(1-(1-d)^{p})^{1/p}$ which for small $d$ is roughly $(pd)^{1/p}$ so $\ell_{p}$ is uniformly strictly monotone at a coordinate.
Furthermore, the diameter of the slice $S(e_{i}^{*},\delta)$ does not depend on the rate at which $\mathfrak{L}_{n}=\|\sum_{i=1}^{n}e_{i}\|$ grows. Indeed, for $h_{M}$ in Example 4.7 we have $\mathfrak{L}_{n}\approx\ln(n)$, for $\ell_{p}$ we have $\mathfrak{L}_{n}=n^{1/p}$, while for the predual $d(w,1)_{*}$ where $w_{n}=1/n$ we have $\mathfrak{L}_{n}\approx n/\ln(n)$.
5. Strictly convex renormings with the LD2P
In the first part of the paper, our aim was to find properties on strictly convex solid Banach lattices of real functions, that prevented such spaces from having the LD2P. In particular, we found that any strictly convex norm on $c_{0}(\Gamma)$ with a 1-symmetric basis, fails the LD2P. One of the main results in this section, tells us that this is close to optimal, even in the non-separable case, since $c_{0}(\Gamma),$ for any set $\Gamma,$ will be shown to admit a strictly convex renorming with a 1-unconditional basis and the LD2P. In the separable case $c_{0}(\mathbb{N})$ with the Nakano norm is such a renorming. The Nakano renorming of $c_{0}(\mathbb{N})$ will play a crucial role in the proofs below and it will for example allow us to construct strictly convex renormings with the LD2P of any Banach space containing complemented copies of $c_{0}(\mathbb{N}).$
We start by a general construction of a norm. Let $\varphi:\ell_{\infty}(\mathbb{N})\to[0,\infty]$ be the function given by $\varphi(x)=\sum_{n=1}^{\infty}x_{n}^{2n}$ where $x=(x_{n}).$ Let $(X,\|\cdot\|)$ be any Banach space. Define the function $\Phi:X^{**}\oplus_{\infty}\ell_{\infty}(\mathbb{N})\to[0,\infty]$ by
(4) $\Phi(z)=\|x\|+\varphi(y),$
where $z=(x,y)$ and $\|\cdot\|$ denotes the canonical norm in the bidual of $X.$ To avoid heavy use of parentheses, we will sometimes just write $\Phi(x,y)$ instead of $\Phi((x,y)),$ $\|\cdot\|$ instead of $\|(\cdot,\cdot)\|$ when $\|\cdot\|$ is a norm on $X^{**}\oplus\ell_{\infty}(\mathbb{N}),$ and $Px$ instead of $P(x)$ when $P$ is an operator on $X^{**}.$ For
$Z_{\infty}$ $:=(X\oplus_{\infty}c_{0}(\mathbb{N}),\|(\cdot,\cdot)\|_{\infty}),$
we have
$Z_{\infty}^{**}$ $=(X^{**}\oplus_{\infty}\ell_{\infty}(\mathbb{N}),\|(\cdot,\cdot)\|_{\infty}).$
We start by collecting some properties of $\Phi.$
Lemma 5.1.
Let $\Phi$ be the function given in (4). We have:
  1. (i)
    The set
    $C=\{(x,y)\in X^{**}\oplus_{\infty}\ell_{\infty}(\mathbb{N}):\Phi(x,y)\leq 1\},$
    is convex, symmetric, and absorbing with $\frac{1}{2}B_{Z_{\infty}^{**}}\subseteq C\subseteq B_{Z_{\infty}^{**}}.$
    In particular, if $(x,y)\in C,$ then $\|(x,y)\|_{\infty}\leq\Phi(x,y).$
  2. (ii)
[…]
  3. (iii)
    $\Phi$ is continuous on $Z_{\infty}.$
  4. (iv)
    For each $(x,y)\in Z_{\infty}^{**},$ if $\lambda>1,$ then
    $\Phi((x,y)/\lambda)\leq\frac{1}{\lambda}\Phi(x,y).$
  5. (v)
[…]
Proof.
(i). Since $\|\cdot\|$ and $\varphi$ are convex and symmetric, $C$ is convex and symmetric as well.
If $\|(x,y)\|_{\infty}\leq 1/2,$ then
$\Phi(x,y)=\frac{1}{2}+\sum_{n=1}^{\infty}\frac{1}{2^{2n}}=\frac{1}{2}+\frac{1}{3}<1,$
so $\frac{1}{2}B_{Z_{\infty}^{**}}\subseteq C$ from which we can also infer that $C$ is absorbing.
If $(x,y)\in\lambda C,$ then
$\frac{1}{\lambda}\|x\|\leq\Phi((x,y)/\lambda)\leq 1,$
so $\|x\|\leq\lambda.$ Similarly,
$\frac{1}{\lambda^{2n}}|y_{n}|^{2n}\leq\varphi(y/\lambda)\leq\Phi((x,y)/\lambda)\leq 1$
and $|y_{n}|\leq\lambda$ for all $n\in\mathbb{N}.$ In particular, if $\lambda=1,$ we get $\|(x,y)\|_{\infty}\leq\Phi(x,y).$
(ii). The Minkowski functional on $C$ is given by
[…]
which means that $(x,y)/\lambda\in C.$ The rest follows from (i).
(iii). We have that
$(x,y)\to x\to\|x\|$
is continuous and that for any NN
$(x,y)\to y\to\sum_{n=1}^{N}y_{n}^{2n}$
is continuous. It follows that $\Phi(\cdot,\cdot)$ is continuous on a dense subset of $Z_{\infty}$ which is enough.
(iv). We have
$\Phi((x,y)/\lambda)=\frac{1}{\lambda}\|x\|+\sum_{n=1}^{\infty}(y_{n}/\lambda)^{2n}\leq\frac{1}{\lambda}\Phi(x,y).$
[…]
$\Phi(z)=\lim_{\lambda\to 1^{+}}\Phi(z/\lambda)\leq 1.$
Assume for contradiction that $\Phi(z)<1$.
[…]
$z^{*}(z)=x^{*}(x)+y^{*}(y)=1.$
[…]
Consider first the case where $y^{*}(y)>0.$ We can write $y=(y_{n})_{n=1}^{\infty}$ and $y^{*}=(y_{n}^{*})_{n=1}^{\infty}$. Note that $|y_{n}|<1$ for all $n\in\mathbb{N}$. Choose $k\in\mathbb{N}$ such that $y_{k}^{*}\cdot y_{k}\neq 0$ and find $t>|y_{k}|$ with
$\Phi(z)+t^{2k}-(y_{k})^{2k}\leq 1.$
Define $w=(x,(v_{n})_{n=1}^{\infty})$ where $v_{n}=y_{n}$ for $n\neq k$ and $v_{k}=t\cdot\operatorname{sign}(y_{k}^{*})$. Since
$\Phi(w)=\Phi(z)+t^{2k}-(y_{k})^{2k}\leq 1,$
[…]
$z^{*}(w)-z^{*}(z)=|y_{k}^{*}|\cdot t-y_{k}^{*}\cdot y_{k}>0.$
So $1\geq z^{*}(w)>z^{*}(z)=1$, which is a contradiction.
Consider now the case $x^{*}(x)>0.$ Choose $t>1$ such that
$\Phi(z)+(t-1)\|x\|\leq 1.$
Define $w=(tx,y).$ Since
$\Phi(w)=\Phi(z)+(t-1)\|x\|\leq 1,$
[…]
$z^{*}(w)-z^{*}(z)=tx^{*}(x)+y^{*}(y)-(x^{*}(x)+y^{*}(y))=(t-1)x^{*}(x)>0.$
So $1\geq z^{*}(w)>z^{*}(z)=1$, which is a contradiction.
We conclude that $\Phi(x,y)=1$ as desired. ∎
Assume for the moment that our $X$ is a strictly convex Banach space and has a normalized shrinking 1-unconditional basis $(e_{\gamma})_{\gamma\in\Gamma}.$ Note that for $e_{\gamma}\in X$ and $e_{n}\in c_{0}(\mathbb{N}),$ we have
[…]
Let $\mathcal{A}=\Gamma\cup\mathbb{N}$ and define by $f_{\alpha}=(e_{\gamma},0)$ when $\alpha=\gamma\in\Gamma$ and $f_{\alpha}=(0,e_{n})$ when $\alpha=n\in\mathbb{N}.$ Since $(e_{n})_{n\in\mathbb{N}}$ is a shrinking 1-unconditional basis for $c_{0}(\mathbb{N}),$ we get by a standard argument that $(f_{\alpha})_{\alpha\in\mathcal{A}}$ is a shrinking 1-unconditional basis for $Z_{\infty}.$ Note that if $(x,y),(u,v)\in Z_{\infty}^{**}$ and $|(x,y)|\leq|(u,v)|$, meaning $|x_{\gamma}|\leq|u_{\gamma}|$ for all $\gamma\in\Gamma$ and $|y_{n}|\leq|v_{n}|$ for all $n\in\mathbb{N}$, then $\Phi(x,y)\leq\Phi(u,v)$.
In the presence of a 1-unconditional basis $\Phi$ will have some additional properties.
Lemma 5.2.
If $X$ is a strictly convex Banach space with a normalized shrinking 1-unconditional basis $(e_{\gamma})_{\gamma\in\Gamma},$ then the function $\Phi$ given in (4) satisfies the following properties:
  1. (i)
    For each $(x,y)\in Z_{\infty}^{**},$ we have
    $\Phi(x,y)=\sup_{G}\Phi(P_{G}^{**}(x,y)),$
    where $P_{G}:Z_{\infty}^{**}\to Z_{\infty}^{**}$ denotes the natural projection onto the finite dimensional subspace $\operatorname{span}(f_{\alpha})_{\alpha\in G}\subset Z_{\infty}^{**}$ and where $G$ is a finite subset of $\mathcal{A}.$
  2. (ii)
[…]
Proof.
(i). Let $\mathcal{F}=\{F\subset\Gamma:|F|<\infty\}$ and $\mathcal{G}=\{G\in\mathcal{A}:|G|<\infty\}$. Order $\mathcal{F}$ by set inclusion and $\mathcal{G}$ by the rule $F_{1}\times\{1,\ldots N_{1}\}\leq F_{2}\times\{1,\ldots N_{2}\}$ if $F_{1}\subseteq F_{2}$ and $N_{1}\leq N_{2}.$ Since $(e_{\gamma}^{*})_{\gamma\in\Gamma}$ is a shrinking 1-unconditional basis for $Z_{\infty}$ and $P_{G}^{**}(x,y)=(P_{F}^{**}x,P_{N}^{**}y)$ for every $G=F\times\{1,\ldots,N\}\subset\mathcal{A},$ we get using Proposition 2.1 that
$\Phi(x,y)$ $=\|x\|+\varphi(y)$
$=\sup_{F\in\mathcal{F}}\|P_{F}^{**}x\|+\sup_{N\in\mathbb{N}}\varphi(P_{N}^{**}y)$
$=\lim_{\mathcal{F}}\|P_{F}^{**}x\|+\lim_{N\to\infty}\varphi(P_{N}^{**}y)$
$=\lim_{\mathcal{G}}\Phi(P_{G}^{**}(x,y))=\sup_{G}\Phi(P_{G}^{**}(x,y)).$
(ii). Using (i) and Lemma 5.1 (iii) we have
$\Phi(x,y)$ $=\sup_{G}\Phi(P_{G}^{**}(x,y))=\sup_{G}\inf_{\lambda>1}\Phi(P_{G}^{**}((x,y)/\lambda))$
$\leq\sup_{G}\inf_{\lambda>1}\Phi((x,y)/\lambda)\leq 1.$
Note that, for the second equality, we used $P_{G}^{**}(x,y)\in Z_{\infty}$, which holds by equation (1) in Proposition 2.1. ∎
Theorem 5.3.
Let $X$ be a strictly convex Banach space and let
$Z:=(X\oplus_{\infty}c_{0}(\mathbb{N}),{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$
be the renorming of $Z_{\infty}$ defined above.
The following statements are true:
  1. (i)
    The Banach space $Z$ is strictly convex.
  2. (ii)
    The Banach space $Z$ is almost square.
  3. (iii)
    The Banach space $Z$ has the LD2P.
If, in addition, $X$ has a shrinking 1-unconditional basis, then the following statements are true:
  1. (iv)
    The Banach space $Z$ has a 1-unconditional basis.
  2. (v)
[…]
Proof.
[…]
$\Phi(a)=\Phi(b)=\Phi\left(\frac{1}{2}(a+b)\right)=1.$
Thus
$\frac{\Phi\left(\frac{1}{2}(a+b)+\frac{1}{2}(a-b)\right)+\Phi\left(\frac{1}{2}(a+b)-\frac{1}{2}(a-b)\right)}{2}$
$=\frac{\Phi(a)+\Phi(b)}{2}=\Phi\left(\frac{1}{2}(a+b)\right)$
It is enough to show that $\frac{1}{2}(a-b)=0$.
To this end it is enough to show that for all $z,w\in Z$ we have that
$\frac{\Phi(z+w)+\Phi(z-w)}{2}>\Phi(z)$
whenever $w\neq 0$.
Let $z=(x,y),w=(u,v)\in Z$ and assume $w\neq 0$. We always have
$\frac{1}{2}\left[\Phi(z+w)+\Phi(z-w)\right]-\Phi(z)$
$=\left[\frac{\|x+u\|+\|x-u\|}{2}-\|x\|\right]+\sum_{n=1}^{\infty}\left[\frac{(y_{n}+v_{n})^{2n}+(y_{n}-v_{n})^{2n}}{2}-y_{n}^{2n}\right]$
$\geq 0$
because the functions $s\mapsto\|s\|$ and $f_{n}(t)=t^{2n},n\in\mathbb{N},$ are convex, hence all expressions in the brackets are non-negative. If $w\neq 0,$ we either have $u\neq 0,$ and thus
$\frac{\|x+u\|+\|x-u\|}{2}-\|x\|>0$
since $\|\cdot\|$ is strictly convex on $X,$ or we have that there exists $n\in\mathbb{N}$ such that $v_{n}\neq 0,$ and thus
$\frac{(y_{n}+v_{n})^{2n}+(y_{n}-v_{n})^{2n}}{2}-y_{n}^{2n}>0$
since $f_{n}$ is strictly convex. In either case we get
$\frac{\Phi(z+w)+\Phi(z-w)}{2}-\Phi(z)>0,$
as desired.
(ii). Put $\Delta=\operatorname{span}\{e_{\gamma}\}_{\gamma\in\Gamma}\oplus_{\infty}c_{00}(\mathbb{N}).$ Since $\Delta$ is dense in $Z,$ it suffices to prove that for $z_{1},\ldots z_{n}\in\Delta\cap S_{Z}$ and $\varepsilon>0,$ there exists $h\in S_{Z}$ such that $\|z_{i}+h\|<1+\varepsilon$ for every $i=1,\ldots,n.$ To this end write $z_{i}=(x_{i},y_{i})$ for $i\in\{1,\ldots,n\}.$ Let $1-\varepsilon/2<r<1$ and find $N>\max\{\operatorname{supp}y_{i}:i=1,\ldots,n\}$ such that $\varphi(y_{i}+re_{N})<\varphi(y_{i})+\varepsilon/2$ for every $i=1,\ldots,n.$ Then for $g=(0,re_{N})$ we get
$\Phi(z_{i}+g)=\|x_{i}\|+\varphi(y_{i}+re_{N})\leq\|x_{i}\|+\varphi(y_{i})+\varepsilon/2=1+\varepsilon/2,$
[…]
which is want we wanted.
(iii). Almost squareness is known to imply the LD2P (see e.g [ALL16, Theorem 1.3 and Proposition 2.5]). Hence (iii) follows from (ii).
[…]
$u_{b}=\sum_{\gamma\in F}b_{\gamma}f_{\eta}\quad\mbox{and}\quad v_{b}=\sum_{n\in E}b_{n}f_{n}$
and
$u_{a}=\sum_{\gamma\in F}a_{\gamma}f_{\eta}\quad\mbox{and}\quad v_{a}=\sum_{n\in E}a_{n}f_{n}$
we have
$\Phi((u_{b},v_{b})/\lambda)\leq\Phi((u_{a},v_{a})/\lambda)$
[…]
Let $G\subset\mathcal{A}$ with $|G|<\infty$. We have
$\|P_{G}z_{\beta}-P_{G}^{**}z\|_{\infty}\to_{\beta}0,$
where $P_{G}$ denotes the projection onto the finite dimensional space $\operatorname{span}(f_{\alpha})_{\alpha\in G}\subset Z$. From Lemma 5.1 (iii) we get
$\Phi(P_{G}z_{\beta})\to_{\beta}\Phi(P_{G}^{**}z).$
By Lemma 5.2 (ii) (and 1-unconditionality of the basis) we have
$\Phi(P_{G}z_{\beta})\leq 1\ \mbox{so that}\ \Phi(P_{G}^{**}z)\leq 1.$
Since $G\subset\mathcal{A}$ with $|G|<\infty$ was arbitrary we get
$\Phi(z)=\sup_{G}\Phi(P_{G}^{**}z)\leq 1$
[…]
The following result should be compared with the main result of [CH24] saying that a Banach space which admits a smooth norm and contains a complemented copy of $\ell_{1}(\mathbb{N})$ has an equivalent norm which is simultaneously smooth and octahedral.
Corollary 5.4.
If a Banach space is strictly convex renormable and contains a complemented copy of $c_{0}(\mathbb{N})$, then it has an equivalent norm which is both strictly convex and almost square.
In particular, any separable Banach space $X$ which contains a copy of $c_{0},$ admits a norm which is simultaneously strictly convex and almost square.
Proof.
A Banach space $Y$ with a strictly convex norm containing a complemented copy of $c_{0}(\mathbb{N})$ is isomorphic to
$Z_{\infty}=X\oplus_{\infty}c_{0}$
where $X$ is a subspace of $Y$. Since the norm of $Y$ is strictly convex the same holds for the norm of $X$. The result now follows from Theorem 5.3.
The particular case follows from the first part and Sobczyk’s lemma. ∎
Corollary 5.5.
For any infinite set $\Gamma$, the Banach space $c_{0}(\Gamma)$ has a strictly convex renorming which is almost square and has a 1-unconditional basis.
Proof.
The result follows from Theorem 5.3 if we just note that $c_{0}(\Gamma)=c_{0}(\Gamma)\oplus_{\infty}c_{0}$ and that we can equip $c_{0}(\Gamma)$ with a strictly convex norm (e.g. Day’s norm [GMZ22, Chapter 9.4.1]). ∎
As mentioned earlier, the predual $d(w,1)_{*}$ of the Lorentz space $d(w,1)$ is never strictly convex. However, we do have the following result.
Corollary 5.6.
The predual $d(w,1)_{*}$ of the Lorentz space $d(w,1),$ admits a strictly convex renorming with a 1-unconditional basis and the LD2P.
Proof.
Since $d(w,1)_{*}$ is separable with a 1-symmetric basis, we get from [HZ95, Proposition 4] that this space has an equivalent 1-symmetric strictly convex renorming $Y$. Since $d(w,1)_{*}$ is non-reflexive and M-embedded, $Y$ contains a complemented copy of $c_{0}(\mathbb{N})$ by [HWW93, Corollary 4.7 (d)]. Hence, denoting this complement by $X,$ which as a subspace of $Y$ is strictly convex, Theorem 5.3 applies to conclude that $d(w,1)_{*}$ admits a strictly convex renorming with a 1-unconditional basis and the LD2P. ∎
Let us end the paper with some questions.
Question 5.7.
Does the predual $d(w,1)_{*}$ of the Lorentz space $d(w,1)$ admit a strictly convex renorming with the LD2P for which the canonical basis of $d(w,1)_{*}$ is 1-symmetric?
Question 5.8.
Does there exist a strictly convex renorming of $\ell_{\infty}(\mathbb{N})$ with the LD2P?
In [AHN+16, Proposition A.1] it was shown that $c_{0}(\mathbb{N})$ with the Nakano norm is not only strictly convex with the LD2P, it is even M-embedded and its dual is smooth. For $\Gamma$ uncountable $\ell_{1}(\Gamma)$ does not have an equivalent smooth norm so such a renorming is not possible for $c_{0}(\Gamma)$. However, Corollary 5.5 gives a strictly convex renorming of $c_{0}(\Gamma)$ which is almost square and has a 1-unconditional basis and we do not know if this can be improved to be M-embedded.
Question 5.9.
Let $\Gamma$ be uncountable. Does there exist an M-embedded strictly convex renorming of $c_{0}(\Gamma)$ with a 1-unconditional basis?
Acknowledgments
Some of this research was carried out when the first author visited the Faculty of Electrical Engineering at the Czech Technical University in Prague in the spring of 2024. He would like to thank his host for the warm hospitality extended during his stay.
The second author was supported by GA23-04776S and SGS22/053/OHK3/1T/13 of CTU in Prague.